# Depth Any Video with Scalable Synthetic Data

### v1

@@ Abstract
Video depth estimation has long been hindered by the scarcity of consistent and scalable ground truth data, leading to inconsistent and unreliable results. In this paper, we introduce Depth Any Video, a model that tackles the challenge through two key innovations. First, we develop a scalable synthetic data pipeline, capturing real-time video depth data from diverse synthetic environments, yielding 40,000 video clips of 5-second duration, each with precise depth annotations. Second, we leverage the powerful priors of generative video diffusion models to handle real-world videos effectively, integrating advanced techniques such as rotary position encoding and flow matching to further enhance flexibility and efficiency. Unlike previous models, which are limited to fixed-length video sequences, our approach introduces a novel mixed-duration training strategy that handles videos of varying lengths and performs robustly across different frame rates—even on single frames. At inference, we propose a depth interpolation method that enables our model to infer high-resolution video depth across sequences of up to 150 frames. Our model outperforms all previous generative depth models in terms of spatial accuracy and temporal consistency. Equal contribution. Corresponding author.
@@ 1 Introduction
To tackle the challenge, in this paper, we propose a solution from two complementary perspectives: (1) constructing a large-scale synthetic video depth dataset and (2) designing a novel framework that leverages powerful visual priors of generative models to effectively handle various real-world videos.
Synthetic Data: Modern video games offer highly realistic graphics and simulate diverse real-world scenarios. For example, car racing games realistically replicate driving environments, while open-world games simulate various complex scenes. Given that modern rendering pipelines often include depth buffers, it becomes possible to extract large-scale, highly accurate video depth data from synthetic environments, which is scalable and cost-effective. In light of this, we construct DA-V, a synthetic dataset comprising 40,000 video clips. DA-V captures a wide range of scenarios, covering various lighting conditions, dynamic camera movements, and intricate object interactions in both indoor and outdoor environments, providing the opportunity for models to generalize effectively to real-world environments.
@@
  • We introduce a large-scale synthetic dataset of 40,000 video depth clips collected from diverse rendering engines. This is the first dataset to leverage a variety of high-fidelity synthetic environments and to verify its generalization capability in real-world scenarios.
@@ 2 Synthetic Data Workflow
Real-time Data Collection. To address the challenges of depth data, we collect a large-scale synthetic dataset comprising approximately 40,000 video clips. A significant portion of this dataset is derived from state-of-the-art synthetic engines, leveraging their ability to generate photorealistic environments with accurate depth information. We extract depth data from a diverse set of virtual environments, carefully selected to encompass a wide range of scene conditions. In Table 1, we compare our dataset with previous public synthetic datasets. To the best of our knowledge, ours is the largest synthetic video dataset covering a wide range of realistic scenes.
@@
Data Filtering. After collecting initial synthetic video data, occasional misalignments between the image and depth are observed. To filter these frames, we first employ a scene cut method [^1] https://github.com/Breakthrough/PySceneDetect to detect scene transitions based on significant color changes. Then, the depth model (detailed in Sec. 3.1), trained on a hand-picked subset of the data, is used to filter out splited video sequences with low depth metric scores. However, this straightforward approach can lead to excessive filtering of unseen data. Therefore, we further use a CLIP (Radford et al., 2021) model to compute semantic similarity between the actual and predicted depth, both colorized from a single channel to three. Finally, we uniformly sample 10 frames from each video segment. If both the median semantic and depth metric scores fall below predefined thresholds, the segment is removed.
@@ 4.3 Zero-shot Depth Estimation
Visual Prior Flow Matching Synthetic Data Training (s) (hours) (%) AbsRel $\downarrow$ $\delta 1\uparrow$
@@
Qualitative Comparisons. Figure 4 presents qualitative monocular depth estimation results across different datasets. It highlights the ability of our model to capture fine-grained details compared to Depth Anything V2 (Yang et al., 2024b), such as the cup in the NYUv2 dataset and the ladder in the ETH3D dataset. Moreover, our model handles objects with similar colors more effectively. For example, on the KITTI dataset, the person’s head blends with the background, which Depth Anything V2 fails to predict. Compared to generative methods like Marigold and ChronoDepth, our model generalizes well to in-the-wild data, offering a better distinction between the sky and foreground. The diversity of DA-V significantly contributes to this, enabling our model to generalize effectively across various environments, particularly in complex, real-world scenarios. Figure 5 further demonstrates qualitative results of depth estimation on open-world videos, covering a wide range of scenarios such as dust and sand, animals, architecture, and human motion presented in both generated and real-world videos. Following Wang et al. (2023), we visualize the changes in estimated depth values over time at the vertical red lines by slicing along the time axis to better capture temporal consistency. Marigold exhibits zigzag artifacts on a per-frame basis, while ChronoDepth displays similar issues at a per-window level. DepthCrafter, with its large overlap between windows and interpolation of overlap space, achieves smoother transitions across windows but still struggles with window-wise flickering. In contrast, our method ensures global consistency by predicting key frames and interpolating intermediate frames, significantly reducing flicker artifacts between windows.
@@ 4.4 Ablation Studies
Synthetic Data. In Table 3, the third and fourth rows show that our collected synthetic data brings gains of 0.7 in $\delta 1$ and 0.4 in AbsRel. The accuracy improvement in outdoor scenes is particularly significant, as shown in Table 2, likely due to the diverse range of outdoor environments in our dataset. Additionally, Figure 4 and 5 demonstrate that models trained on our synthetic data generalize well to in-the-wild scenarios, showcasing both the realism and effectiveness of our synthetic data.
@@ 6 Conclusion
We present Depth Any Video, a novel approach for versatile image and video depth estimation, powered by generative video diffusion models. Leveraging diverse and high-quality depth data collected from diverse synthetic environments, our model could generate temporally consistent depth sequences with fine-grained details across a broad spectrum of unseen scenarios. Equipped with a mixed-duration training strategy and frame interpolation, it generalizes effectively to videos of various lengths and resolutions. Compared to previous generative depth estimation models, our approach sets a new state-of-the-art in performance while significantly enhancing efficiency.

### v2

@@ Abstract
Video depth estimation has long been hindered by the scarcity of consistent and scalable ground truth data, leading to inconsistent and unreliable results. In this paper, we introduce Depth Any Video, a model that tackles the challenge through two key innovations. First, we develop a scalable synthetic data pipeline, capturing real-time video depth data from diverse virtual environments, yielding 40,000 video clips of 5-second duration, each with precise depth annotations. Second, we leverage the powerful priors of generative video diffusion models to handle real-world videos effectively, integrating advanced techniques such as rotary position encoding and flow matching to further enhance flexibility and efficiency. Unlike previous models, which are limited to fixed-length video sequences, our approach introduces a novel mixed-duration training strategy that handles videos of varying lengths and performs robustly across different frame rates—even on single frames. At inference, we propose a depth interpolation method that enables our model to infer high-resolution video depth across sequences of up to 150 frames. Our model outperforms all previous generative depth models in terms of spatial accuracy and temporal consistency. The code and model weights are open-sourced. Equal Contribution. This work was done during his internship at Shanghai AI Laboratory.
@@ 1 Introduction
To tackle the challenge, we propose a solution from two complementary perspectives: (1) demonstrating a scalable data collection pipeline to expand data and (2) designing a novel framework that leverages powerful visual priors of generative models to effectively handle various real-world videos.
Synthetic Data: Modern virtual environments offer highly realistic graphics and simulate diverse real-world scenarios. For example, driving simulations accurately recreate real-world road conditions, while open-ended virtual worlds depict various complex scenes. Given that modern graphics rendering pipelines often incorporate depth buffers, real-time rendering engines (e.g., Unreal Engine, Unity) offer functionalities for generating and accessing high-quality video depth data, enabling its extraction during interactive sessions. In light of this, we construct DA-V, a synthetic dataset comprising 40,000 video clips collected from diverse interactive virtual environments over two weeks by 10 human players. DA-V captures a wide range of scenarios, covering various lighting conditions, dynamic camera movements, and intricate object interactions in both indoor and outdoor environments, providing the opportunity for models to generalize effectively to real-world environments.
@@
  • We demonstrate a feasible path for scaling depth data from diverse virtual environments and systematically validate that high-fidelity synthetic data can improve the generalization of video depth models to real-world scenarios.
@@ 2 Synthetic Data Workflow
Rendering Pipeline Overview. Modern rendering pipelines are central to transforming geometric data into visual output, which involves three key stages: 1) Vertex Processing: Geometric data is first processed to determine the position and attributes of each vertex in 3D space. 2) Rasterization: The processed geometric data is then converted into pixels. This stage determines which pixels on the screen correspond to parts of the 3D models, establishing depth and color data for each pixel. 3) Shading: Shaders compute the lighting, coloring, and textures. Deferred shading, a popular technique, separates lighting calculations from geometry rendering to optimize processing. Buffers such as depth buffers store intermediate data essential for accurate lighting and shading.
Real-time Data Collection. Following prior works (Richter et al., 2016; Krähenbühl, 2018), we leverage ReShade (ReShade Contributors, 2024) to access depth information available in the rendering process and OBS (Contributors, 2024) to capture synchronized depth and color data from the screen. The collected dataset encompasses diverse scenes and conditions, including modern virtual environments such as Cities: Skylines II with its expansive urban landscapes and Grand Theft Auto V with its varied open-world settings. Notably, the use and sharing of certain game content are permitted under specific terms and conditions (Rockstar Games, 2024; Paradox Interactive, 2024), including non-commercial use and restrictions on story spoilers. To ensure compliance, we further remove UI elements and on-screen text using a ReShade plugin [^1] https://github.com/4lex4nder/ReshadeEffectShaderToggler and manually filter specific scenes to prevent unintended storyline disclosures.
@@
Data Filtering. After collecting initial synthetic data, occasional misalignments between the image and depth are observed, especially during menu switches. To filter these frames, we first employ a scene cut method [^2] https://github.com/Breakthrough/PySceneDetect to detect scene transitions based on significant color changes. Then, the depth model (detailed in Sec. 3.1), trained on a hand-picked subset of collected data, is used to filter out splited video sequences with low depth metric scores. However, this straightforward approach can lead to excessive filtering of unseen data. We further use a CLIP (Radford et al., 2021) model to compute semantic similarity between the actual and predicted depth, both colorized from a single channel to three. Finally, we uniformly sample 10 frames from each video segment. If both the median semantic and depth metric scores fall below predefined thresholds, the segment is removed. After filtering, we obtain a curated synthetic dataset of 40,000 video clips, each recorded at 30 FPS and lasting 5 seconds, as shown in Table 1.
@@ 4.3 Zero-shot Depth Estimation
Visual Prior Flow Matching Data Training (s) (hours) (%) AbsRel $\downarrow$ $\delta 1\uparrow$
@@
Qualitative Comparisons. Figure 4 presents qualitative monocular depth estimation results across different datasets. It highlights the ability of our model to capture fine-grained details compared to Depth Anything V2 (Yang et al., 2024b), such as the cup in the NYUv2 dataset and the ladder in the ETH3D dataset. Moreover, our model handles objects with similar colors more effectively. For example, on the KITTI dataset, the person’s head blends with the background, which Depth Anything V2 fails to predict. Compared to generative methods like Marigold and ChronoDepth, our model generalizes well to in-the-wild data, offering a better distinction between the sky and foreground. The DA-V dataset significantly contributes to this, enabling our model to generalize effectively across various environments, particularly in complex, real-world scenarios. Figure 5 further demonstrates qualitative results of depth estimation on open-world videos, covering a wide range of scenarios such as dust and sand, animals, architecture, and human motion presented in both generated and real-world videos. Following Wang et al. (2023), we visualize the changes in estimated depth values over time at the vertical red lines by slicing along the time axis to better capture temporal consistency. Marigold exhibits zigzag artifacts on a per-frame basis, while ChronoDepth displays similar issues at a per-window level. DepthCrafter, with its large overlap between windows and interpolation of overlap space, achieves smoother transitions across windows but still struggles with window-wise flickering. In contrast, our method ensures global consistency by predicting key frames and interpolating intermediate frames, significantly reducing flicker artifacts between windows.
@@ 4.4 Ablation Studies
Synthetic Data. In Table 3, the third and fourth rows show that our collected DA-V dataset brings gains of 0.7 in $\delta 1$ and 0.4 in AbsRel. The accuracy improvement in outdoor scenes is particularly significant, as shown in Table 2, likely due to the diverse range of outdoor environments in our dataset. Additionally, Figure 4 and 5 demonstrate that models trained on our synthetic data generalize well to in-the-wild scenarios, showcasing both the realism and effectiveness of our synthetic data.
@@ 6 Conclusion
We present Depth Any Video, a novel approach for versatile image and video depth estimation, powered by generative video diffusion models. Leveraging diverse and high-quality depth data collected from virtual environments, our model could generate temporally consistent depth sequences with fine-grained details across a broad spectrum of unseen scenarios. Equipped with a mixed-duration training strategy and frame interpolation, it generalizes effectively to videos of various lengths and resolutions. Compared to previous generative depth estimation models, our approach sets a new state-of-the-art in performance while significantly enhancing efficiency.